OpenStudy (anonymous):

What is the value of log81 3? Show your work. Can someone help me please?

OpenStudy (welshfella):

the base of the log is 81 so we can write 81^x = 3 find the value of x

OpenStudy (welshfella):

HInt: 3 is the 4th root of 81

OpenStudy (anonymous):

Okay would it be 4

OpenStudy (welshfella):

no its 1/4 because 81 ^ (1/4) = 3 exponent 1/2 = square root exponent 1/4 = 4th root etc

OpenStudy (welshfella):

for fractional exponents:-